# journal: TOG

### v1

@@ 3.2. Network Architecture
The deletion network predicts a mask label for each vertex with a deletion head, which processes the encoded states of every three coordinate tokens corresponding to a vertex. While using the same OPT architecture, the attention layers are switched to perform bi-directional attention for each position to capture the global context. After the inference, we aggregate $\mathcal{V}^{\prime}_{r}$ according to Eq. 2, and then obtain the deleted mesh with Eq. 1.
3.3. Vertex-aligned Speculator
As we perform mesh addition autoregressively, generating an $n$-token sequence requires $n$ forward passes, significantly limiting mesh generation speed for interactive applications. One effective way of accelerating generation is to use a speculator to generate multiple tokens in a single run. In contrast to natural language, a tokenized mesh sequence contains strong low-level structure – each vertex $\boldsymbol{v}$ is defined by exactly $3$ tokens $V_{\{x,y,z\}}$ representing its $x$, $y$, and $z$ coordinates. Aligning the speculator with vertex tokens by letting it predict only $V_{\{y,z\}}$, as shown in Fig. 3, ensures a consistent context of input and output for the speculator.
We adopt the MLP speculator described in (Wertheimer et al., 2024). In natural language processing tasks, a speculator can be trained with a frozen pre-trained large language model to accelerate generation. However, in our experiments, we found that this results in performance degradation. We thus train our speculator jointly with the the mesh addition task, so that the addition transformer learns to inform the speculator with contextual information in hidden states.
@@ 3.3. Vertex-aligned Speculator
(5) $P\left(V^{\prime}_{\{y,z\}}\right)=\text{Speculator}\left(E_{x},V^{\prime}_{x}\right),$
@@
3.4. Self-Supervised Data Generation for Mesh Editing
@@
By decomposing sketch-conditioned mesh editing into separate deletion and addition subtasks, we can generate training data for each subtask from 3D shape datasets without requiring real mesh editing sequences.
@@
Given a complete 3D mesh sample $\mathcal{M}_{c}$, we select its triangles to form two mutually exclusive subsets $\mathcal{M}_{r}$ and $\mathcal{M}_{k}$, representing the edited and unedited parts, respectively. Note that $\mathcal{M}_{r}\cup\mathcal{M}_{k}=\mathcal{M}_{c}$ is not required, as intermediate meshes during editing can be incomplete before deletion or after addition. To achieve this, we sample two volumes $\mathcal{L}$ and $\mathcal{L}_{k}\subsetneq\mathcal{L}$ within the bounding volume of $\mathcal{M}_{c}$. An example is shown in Fig. 4. For each volume, we crop a partial mesh by selecting all triangles containing at least one vertex within the given volume:
(6) =$\mathcal{M}=\left\{\mathcal{F}\in\mathcal{M}_{c}|\exists\boldsymbol{p}\in\mathcal{F}:\boldsymbol{p}\in\mathcal{L}\right\},$
@@
and the same for $\mathcal{M}_{k}$, $\mathcal{L}_{k}$. We then compute $\mathcal{M}_{r}=\mathcal{M}\setminus\mathcal{M}_{k}$, which is mutually exclusive from $\mathcal{M}_{k}$. We provide more details about the volume sampling and mesh cropping in the supplementary. The obtained meshes are then used to supervise both the addition and the deletion network in the tasks defined in Sec. 3.1.
@@
Figure 4. An example of training data generation: the volume $\mathcal{L}$ is sampled to cover a large portion from the top, while $\mathcal{L}_{k}$ covers the chair backrest. The mesh $\mathcal{M}$ and $\mathcal{M}_{k}$ contains triangles with any vertex in $\mathcal{L}$ and $\mathcal{L}_{k}$, respectively. $\mathcal{M}_{r}$ is the difference between $\mathcal{M}$ and $\mathcal{M}_{k}$. The red and black line strokes in sketch $\mathcal{I}$ corresponds to $\mathcal{M}_{k}$, and $\mathcal{M}_{r}$, respectively.
Sketch Generation. We generate synthetic sketches corresponding to the processed mesh data. During training, this is done automatically through Canny edge detection; at test time, any human-drawn sketch can be used. Fig. 4 shows our data generation process.
Automatic sketch generation is done by applying Canny edge detection on the rendered depth and normal images of $\mathcal{M}$. We then merge the two edge detection results to obtain the corresponding synthetic sketch used for training. To get mutually exclusive sets of line strokes $\mathcal{I}_{\{r,k\}}$, which correspond to different parts of the mesh, we further render a visibility mask of $\mathcal{M}_{r}$ in the same camera view and collect all line strokes within the mask as $\mathcal{I}_{r}$. $\mathcal{I}_{k}$ is then defined as the line strokes not in $\mathcal{I}_{r}$. During training, we randomly sample camera views from the upper unit viewing hemisphere, with azimuth angles in $[-90^{\circ},90^{\circ}]$ and elevation angles in $[0^{\circ},60^{\circ}]$.
4. Experiments
4.1. Implementation Details
@@
Data. Following MeshGPT (Siddiqui et al., 2023) and MeshAnythingV2 (Chen et al., 2024e), we preprocess and filter the ShapeNet dataset (Chang et al., 2015) to obtain approximately $28$k meshes from $55$ categories, which each have a face count $<768$ for training. We sample around $500$ meshes for validation and $1000$ meshes for testing. Other meshes are used for training.
@@
Training. Both the mesh addition and deletion networks are initialized using the OPT weights from the MeshAnythingV2 checkpoint, with the RADIO image model frozen throughout training. We employ a learning rate schedule from $1\times 10^{-4}$ to $1\times 10^{-6}$. For the addition model, training begins with a batch size of $56$ for $128$ epochs, followed by $20$ epochs with a batch size of $224$. The deletion model is trained with a batch size of $32$ for $128$ epochs. Using 4 NVIDIA A100 GPUs, the training process takes approximately $5$ days for the addition network and $2$ days for the deletion network.
User Interface. We develop a Gradio (Abid et al., 2019) user interface to demonstrate our method in an interactive editing environment, accessible from browsers on all types of devices (PC, iPad, etc.). As visualized in Fig. 1, the interface contains a sketchpad for users to draw (addition) and erase (deletion) line strokes and a three.js (Cabello and Contributors, 2010) viewer displaying the mesh in real-time. Users can submit their sketches to the addition or deletion job and can always re-sketch and re-submit the job when the outcome should be iterated on. When an edit is accepted, the interface refreshes the sketch in the sketchpad to match the current mesh for further edit, which corresponds to the process depicted in Fig. 2. We refer to our supplementary video for live demos.
4.2. Baselines and Metrics
@@
In our experiments, we evaluate the generated mesh quality as well as its consistency with the input sketch. For mesh quality, we report the Chamfer Distance (CD) to the target mesh. A smaller CD indicates better geometric consistency with the target. Additionally, for perceptual mesh quality evaluation we employ shading-image-based Fréchet Inception distance (FID) following MeshGPT (Siddiqui et al., 2023) and SENS (Binninger et al., 2024). We also evaluate sketch-based CLIP (Radford et al., 2021) similarity, following LAS (Zheng et al., 2023), to measure the similarity between the generated mesh and the sketch. We additionally introduce sketch-based LPIPS (Zhang et al., 2018) to evaluate sketch-to-mesh correspondence: while CLIP shows the similarity of the global context, LPIPS focuses more on local features. For our method, we analyze runtime with the token generation speed in $tokens/second$ (T/s).
@@
We compare with two state-of-the-art sketch-conditioned 3D generation methods, LAS (Zheng et al., 2023) and SENS (Binninger et al., 2024) as baselines. Since these methods produce signed distance fields or occupancy grids, we use marching cubes to convert their results to meshes. We alternatively use MeshAnythingV2 (Chen et al., 2024e) to post-process LAS and SENS results (LAS-MA and SENS-MA, respectively), producing more artistically-styled meshes for these baselines.
4.3. Experimental Setup
We use the test set described in Sec. 4.1 and automatically generate sketches from one of five views: top-left, top-front-left, top-front, top-front-right, or top-right, for the generation task. To further show the robustness of our method, we use the IKEA dataset (Lim et al., 2013; Sun et al., 2018), which contains $188$ furniture models. Additionally, we randomly select $50$ shapes of airplanes, chairs, and lamps (on which SENS is trained) from the two datasets and manually edit the automatically generated sketches to build an evaluation dataset for sketch-based mesh editing.
As SENS is category-specific and is only trained on airplanes, chairs, and lamps, we evaluate it only on these categories. For the editing task, we let LAS generate from scratch conditioned on the edited sketch since it does not support localized editing; as SENS was designed for editing, and we ran its editing using its original design. For our method, while providing only the original sketch and the edited sketch, we first automatically generate input for deletion by identifying the difference between the two sketches, then perform the deletion-addition operation.
4.4. Comparison to State of the Art
@@
Table 1. Evaluation on sketch-conditioned mesh generation. We compare with LAS (Zheng et al., 2023), SENS (Binninger et al., 2024), and further use MeshAnythingV2 (MA) (Chen et al., 2024e) to post-process LAS and SENS outputs. The unit of CD is 0.001. MeshPad outperforms baselines in both mesh quality and consistency with input sketches.
@@
Table 2. Sketch-based metrics and unary perceptual study ratings (ranged $1$-$5$) on generation and editing results of our hand-drawn sketch evaluation set. We benchmark against LAS (Zheng et al., 2023) and SENS (Binninger et al., 2024), with outputs further refined using MeshAnythingV2 (MA) (Chen et al., 2024e). For mesh generation, participants evaluate mesh quality (GQ) and sketch matching (GM). For mesh editing, participants rate edited mesh quality (EQ), edited sketch matching (EM), and edited mesh consistency (EC) of the unedited part.
@@
Table 3. Binary perceptual study. Participants choose their preferred output between our method and each baseline (LAS (Zheng et al., 2023), SENS (Binninger et al., 2024), and their results post-processed by MeshAnythingV2 (MA) (Chen et al., 2024e)). Our method is preferred by a large margin for both generation and editing.
@@
Mesh Editing. Based on the results of the perceptual study, together with the sketch metrics reported in Tabs. 2 and 3, our method achieves the best LPIPS and CLIP scores among baselines and reaches the highest rating and preference among users. From the visual comparisons provided in Fig. 7, we notice that baseline methods suffer not only from poor editing results but also from inconsistency in unedited regions. This is because these methods regenerate the whole shape after the edit, can change regions not intended to be edited. Our method solves this problem by decoupling the editing procedure into deletion and addition, ensuring that only edited regions are changed.
Figure 5. Novelty of generated shapes. We retrieve the nearest neighbors of the generated meshes by first aligning them to each train mesh with Iterative Closest Point (ICP) and then computing the Chamfer distance. Our method can generate new shapes different from those in the train set.
Novelty Analysis. Fig. 5 shows a qualitative novelty analysis, retrieving for our generated meshes the 3 nearest neighbors from the training set, based on Chamfer distance. Our generated shapes differ from their nearest neighbors in both overall shape and triangulation, which demonstrates that our method can generate new shapes.
Figure 6. Comparisons between our interactive generation and editing vs. single-step generation. Left: meshes interactively created using our method, which effectively uses multiple localized generation steps to achieve high complexity. Middle: synthetic sketches generated from the final meshes via Canny Edge detection. Right: meshes generated by applying one pass of the addition network (without iterative editing) on the middle sketches. Our interactive editing pipeline enables creating intricate shapes that can be challenging to generate in one step.
4.5. Ablations
@@
Joint training provides context-rich hidden states as speculator input. We find that the vanilla speculator training scheme with frozen transformer (Wertheimer et al., 2024) does not work in our mesh generation setting, as the hidden states from our pre-trained transformer do not contain enough contextual information. As shown in Tab. 4, by instead training the speculator jointly with the OPT, our model achieves strong mesh quality performance. In contrast, training the speculator with the OPT frozen results in poor performance. This indicates the OPT learns to encode contexts into the hidden states when jointly trained.
Vertex-alignment preserves mesh quality while achieving comparable speedup as vanilla MLP speculator. Comparing no speculator (w/o speculator) and speculator without vertex alignment (w/o vert-alignment), we can see that the speculator, without aligning to vertex tokens, results in a notable speedup, but at large penalty cost in mesh quality and sketch-to-mesh correspondence. By aligning the speculator to vertex tokens, our method reaches the same level of sketch-to-mesh correspondence as the no speculator version while maintaining the naive speculator’s advantage in generation speed. Additionally, applying the vertex-aligned speculator results in slightly higher metrics in mesh quality. This is likely due to the vertex-aligned speculator reducing the complexity of the linear prediction head. Rather than predicting control tokens and all $x$, $y$, and $z$ coordinate tokens, it is trained to predict only control tokens and $x$ coordinate tokens, which reduces the overall prediction error. Thus, equipped with the vertex-aligned speculator, our method achieves strong performance and fast generation speeds (approximately $26.7$ faces per second considering the average face density of the mesh sequences).
Runtime. Our mesh addition, which typically generates $\approx\nicefrac{{1}}{{5}}$ of a shape, runs in about 1 to 5 seconds (depending on the number of faces added), enabling interactive runtimes for editing.
Limitations. While our method produces interactive editing for 3D meshes, our transformer-based approach is bound by the memory size to a maximum sequence length, resulting in a limited number of mesh triangles ($768$ triangles in our experiments) that can be processed. Although there are methods (Hao et al., 2024; Tang et al., 2024) that extend the limit of the token amount in mesh generation, the upper bound (thousands of triangles) is still far below the requirement of modern computer graphics applications (millions of triangles). Thus, we demonstrate the efficacy of our approach on isolated 3D shapes, but would require paradigms in learned mesh representations in order to generate and edit larger-scale scenes.
Table 4. Ablation study. We ablate our speculator, vertex alignment of our speculator (using a common MLP speculator (Wertheimer et al., 2024) with the same length of $2$), and its joint training with our mesh transformer. The unit of CD is 0.001. T/s represents tokens per second and is measured on NVIDIA A100. Our vertex-aligned speculator accelerates generation by 2.16×2.16$\times$2.16 × without loss in mesh quality. Compared to an MLP speculator without vertex-alignment, which suffers from quality loss, our method runs at a comparable speed.
@@
5. Conclusion

### v2

@@ 3.2 Network Architecture
The deletion network predicts a mask label for each vertex with a deletion head, which processes the encoded states of every three coordinate tokens corresponding to a vertex. While using the same OPT architecture, the deletion network does not share weights with the addition network. The attention layers in the deletion network are switched to perform bi-directional attention for each position to capture the global context. After the inference, we aggregate $\mathcal{V}^{\prime}_{r}$ according to Eq. 2, and then obtain the deleted mesh with Eq. 1.
3.3 Vertex-Aligned Speculator
As we perform mesh addition autoregressively, generating an $n$-token sequence requires $n$ forward passes, significantly limiting mesh generation speed for interactive applications. One effective way of accelerating generation is to use a speculator to generate multiple tokens in a single run. In contrast to natural language, a tokenized mesh sequence contains strong low-level structure—each vertex $\boldsymbol{v}$ is defined by exactly $3$ tokens $V_{\{x,y,z\}}$ representing its $x$, $y$, and $z$ coordinates. Aligning the speculator with vertex tokens by letting it predict only $V_{\{y,z\}}$, as shown in Fig. 3, ensures a consistent context of input and output for the speculator.
We adopt the MLP speculator described in [57]. In natural language processing tasks, a speculator can be trained with a frozen pre-trained large language model to accelerate generation. However, in our experiments, we found that this results in performance degradation. We thus train our speculator jointly with the the mesh addition task, so that the addition transformer learns to inform the speculator with contextual information in hidden states.
@@ 3.3 Vertex-Aligned Speculator
$P\left(V^{\prime}_{\{y,z\}}\right)=\text{Speculator}\left(E_{x},V^{\prime}_{x}\right),$ (5)
@@
3.4 Training Data Generation for Mesh Editing
@@
By decomposing sketch-conditioned mesh editing into separate deletion and addition subtasks, we can generate training data for each subtask from 3D shape datasets without requiring real mesh editing sequences. Fig. 4 shows our data generation process.
@@
Given a complete 3D mesh sample $\mathcal{M}_{c}$, we select its triangles to form two mutually exclusive subsets $\mathcal{M}_{r}$ and $\mathcal{M}_{k}$, representing the edited and unedited parts, respectively. Note that $\mathcal{M}_{r}\cup\mathcal{M}_{k}=\mathcal{M}_{c}$ is not required, as intermediate meshes in interactive editing can be incomplete before deletion or after addition. To achieve this, we sample two volumes $\mathcal{L}$ and $\mathcal{L}_{k}\subsetneq\mathcal{L}$ within the bounding volume of $\mathcal{M}_{c}$. An example is shown in Fig. 4. For each volume, we crop a partial mesh by selecting all triangles containing at least one vertex within the given volume:
=$\mathcal{M}=\left\{\mathcal{F}\in\mathcal{M}_{c}|\exists\boldsymbol{p}\in\mathcal{F}:\boldsymbol{p}\in\mathcal{L}\right\},$ (6)
@@
and the same for $\mathcal{M}_{k}$, $\mathcal{L}_{k}$. We compute $\mathcal{M}_{r}=\mathcal{M}\setminus\mathcal{M}_{k}$, which is mutually exclusive from $\mathcal{M}_{k}$. The volume sampling and mesh cropping details can be found in the supplemental. We use the resulting meshes to supervise both the addition and the deletion networks defined in Sec. 3.1.
@@
Figure 4: An example of training data generation: the volume $\mathcal{L}$ is sampled to cover a large portion from the top, while $\mathcal{L}_{k}$ covers the chair backrest. The mesh $\mathcal{M}$ and $\mathcal{M}_{k}$ contains triangles with any vertex in $\mathcal{L}$ and $\mathcal{L}_{k}$, respectively. $\mathcal{M}_{r}$ is the difference between $\mathcal{M}$ and $\mathcal{M}_{k}$. The red and black line strokes in sketch $\mathcal{I}$ corresponds to $\mathcal{M}_{k}$, and $\mathcal{M}_{r}$, respectively.
Sketch Generation. During training, we generate synthetic sketches corresponding to the processed mesh data by performing Canny edge detection on mesh normal and depth renderings. We randomly sample rendering viewpoints and apply image augmentations to synthesize different line drawing styles. The trained model can thus handle human freehand sketches from various viewpoints. We provide the sketch generation details in the supplemental.
Figure 5: Visual comparison of generation and editing results. We show generation and editing results of LAS [63], SENS [7], and their results post-processed by MeshAnythingV2 (MA) [17]. Our method produces 3D meshes that not only match the sketch inputs and edits, but also offer high fidelity.
4 Experiments
4.1 Implementation Details
@@
Data. Following MeshGPT [51] and MeshAnythingV2 [17], we preprocess and filter the ShapeNet dataset [9] to obtain approximately $28$k meshes from $55$ categories, which each have a face count $<768$ for training. We sample around $500$ meshes for validation and $1000$ meshes for testing. Other meshes are used for training.
@@
Training. Both the mesh addition and deletion networks are initialized using the OPT weights from the MeshAnythingV2 checkpoint. The RADIO model is frozen during training. We employ a learning rate schedule from $1\times 10^{-4}$ to $1\times 10^{-6}$. For the addition model, training begins with a batch size of $56$ for $128$ epochs, followed by $20$ epochs with a batch size of $224$. The deletion model is trained with a batch size of $32$ for $128$ epochs. Using 4 NVIDIA A100 GPUs, the training process takes approximately $5$ days for the addition network and $2$ days for the deletion network.
4.2 Metrics and Baselines
@@
We evaluate the generated mesh quality as well as its consistency with the input sketch. For mesh quality, we report the Chamfer Distance (CD) to the target mesh. A smaller CD indicates better geometric consistency with the target. For perceptual mesh quality evaluation, we employ shading image-based Fréchet Inception distance (FID) following MeshGPT [51] and SENS [7]. We also evaluate sketch-based CLIP [47] similarity, following LAS [63], to measure the similarity between the generated mesh and the sketch. We additionally introduce sketch-based LPIPS [60] to evaluate sketch-to-mesh correspondence: while CLIP shows the similarity of the global context, LPIPS focuses more on local features. For our method, we analyze runtime with the token generation speed in $tokens/second$ (T/s). For this, we measure the time of generating the first $300$ tokens of the addition task, which eliminates the speed difference caused by the varying lengths of generated sequences.
@@
We compare with two state-of-the-art sketch-conditioned 3D generation methods, LAS [63] and SENS [7]. These methods produce signed distance fields or occupancy grids, and we use marching cubes to convert their outputs to triangle meshes. Since LAS and SENS were trained on fewer categories than our method, we report the metrics considering only the results of their respective trained categories (LAS: chair, car, airplane, table, and rifle; SENS: airplane, chair, and lamp). For our method, the metrics are computed on all data in the test set. We alternatively use MeshAnythingV2 [17] to post-process LAS and SENS results (LAS-MA and SENS-MA, respectively), producing more artistically-styled meshes for these baselines.
Figure 6: Sequence visualization of mesh creation with MeshPad. We show the input sketch on the top left of each of our method outputs. Green and purple arrows represent addition and deletion, respectively. The “$+$” inside a green arrow indicates multiple addition steps.
4.3 Experimental Setup
We use the test set described in Sec. 4.1 and automatically generate sketches from one of five views: top-left, top-front-left, top-front, top-front-right, or top-right, for the generation task. To further show the robustness of our method, we use the IKEA dataset [36, 53], which contains $188$ furniture models. We use automatically generated sketches (follow Sec. 3.4) for generation tasks because it helps to analyze the sketch-to-mesh correspondence by the CLIP and LPIPS scores. In our qualitative experiments we show our method generalizes to hand-drawn sketches. Additionally, we randomly select $50$ shapes of airplanes, chairs, and lamps (on which SENS is trained) from the two datasets, automatically generate their sketches for generation, and then manually edit the sketches for one-step editing evaluation.
For the generation task, we perform one-step addition operation for our method. We provide additional camera view conditions in all LAS runs as suggested in [63]. We also apply dilation and resizing to sketches to match the input to the training data of LAS. As SENS is category-specific, we use its checkpoint corresponding to the category of the target mesh for each evaluation run.
For the editing task, we let LAS generate from scratch conditioned on the edited sketch, since it does not support localized editing. SENS was designed for editing, and we ran its editing using its original design. For our method, while providing only the original sketch and the edited sketch, we first automatically generate input for deletion by identifying the difference between the two sketches, then perform the deletion-addition operation.
4.4 Comparison to State of the Art
@@
Table 1: Evaluation on sketch-conditioned mesh generation. We compare with LAS [63], SENS [7], and further use MeshAnythingV2 (MA) [17] to post-process LAS and SENS outputs. The unit of CD is 0.001. Our results are generated by one addition operation from scratch. MeshPad outperforms baselines in both mesh quality and consistency with input sketches.
@@
Table 2: Sketch-based metrics and unary perceptual study ratings (ranged $1$-$5$) on generation and editing results of our hand-drawn sketch evaluation set. We benchmark against LAS [63] and SENS [7], with outputs further refined using MeshAnythingV2 (MA) [17]. For mesh generation, participants evaluate mesh quality (GQ) and sketch matching (GM). For mesh editing, participants rate edited mesh quality (EQ), edited sketch matching (EM), and edited mesh consistency (EC) of the unedited part.
@@
Table 3: Binary perceptual study. Participants choose their preferred output between our method and each baseline (LAS [63], SENS [7], and their results post-processed by MeshAnythingV2 (MA) [17]). Our method is preferred by a large margin for both generation and editing.
Mesh Editing. Based on the results of the perceptual study, together with the sketch metrics reported in Tabs. 2 and 3, our method achieves the best LPIPS and CLIP scores among baselines and reaches the highest rating and preference among users. From the visual comparisons provided in Fig. 5, we notice that baseline methods suffer not only from poor editing results but also from inconsistency in unedited regions. This is because these methods regenerate the whole shape after the edit, can change regions not intended to be edited. Our method solves this problem by decoupling the editing procedure into deletion and addition, ensuring that only edited regions are changed. We present editing sequences with real human inputs in Fig. 6, demonstrating our method’s generalizability. Moreover, our method can edit meshes that are not generated by our networks, which is difficult for baselines, and the results are provided in the supplemental.
Figure 7: Novelty of interactively created shapes. We retrieve the nearest neighbors of the meshes by first aligning them to each train mesh with Iterative Closest Point (ICP) and then computing the Chamfer distance. MeshPad can generate new shapes different from those in the train set and can generate shapes beyond the categories of the training data (third row).
Novelty Analysis. Fig. 7 shows a qualitative novelty analysis, retrieving for our interactively created meshes the 3 nearest neighbors from the training set, based on Chamfer distance. Our generated shapes differ from their nearest neighbors in both overall shape and triangulation, which demonstrates that our method can generate new shapes.
4.5 Ablations
@@
Joint training provides context-rich hidden states as speculator input. We find that the vanilla speculator training scheme with frozen transformer [57] does not work in our mesh generation setting, as the hidden states from our pre-trained transformer do not contain enough contextual information. As shown in Tab. 4, by instead training the speculator jointly with the OPT, our model achieves strong mesh quality performance. In contrast, training the speculator with the OPT frozen results in poor performance. This indicates the OPT learns to encode contexts into the hidden states when jointly trained.
Vertex-alignment preserves mesh quality while achieving comparable speedup as vanilla MLP speculator. Comparing no speculator (w/o speculator) and speculator without vertex alignment (w/o vert-alignment), we can see that the speculator, without aligning to vertex tokens, results in a notable speedup, but at large penalty cost in mesh quality and sketch-to-mesh correspondence. By aligning the speculator to vertex tokens, our method reaches the same level of sketch-to-mesh correspondence as the no speculator version while maintaining the naive speculator’s advantage in generation speed. Additionally, applying the vertex-aligned speculator results in slightly higher metrics in mesh quality. This is likely due to the vertex-aligned speculator reducing the complexity of the linear prediction head. Rather than predicting control tokens and all $x$, $y$, and $z$ coordinate tokens, it is trained to predict only control tokens and $x$ coordinate tokens, which reduces the overall prediction error. Thus, equipped with the vertex-aligned speculator, our method achieves strong performance and fast generation speeds (approximately $26.7$ faces per second considering the average face density of the mesh sequences). As a result, mesh additions, which typically generate about $\nicefrac{{1}}{{5}}$ of a complete shape, run in about 1 to 5 seconds (depending on the number of faces added), enabling interactive runtimes for editing.
Limitations. Similar to recent mesh generation methods [51, 14], our transformer model is limited by context length. Although there are recent attempts [27] to extend the context length, their outputs (thousands of triangles) still fall below the requirement of modern computer graphics applications (millions of triangles). Leveraging techniques, like token compression [30], is a promising avenue to explore. Nevertheless, our approach achieves interactive mesh generation and editing with easy-to-use sketches, enabling iterative creation of novel and complex 3D shapes.
@@
Table 4: Ablation study. We ablate our speculator, vertex alignment of our speculator (using a common MLP speculator [57] with the same length of $2$), and its joint training with our mesh transformer. The unit of CD is 0.001. T/s represents tokens per second and is measured on NVIDIA A100. Our vertex-aligned speculator accelerates generation by 2.16×2.16$\times$2.16 × without loss in mesh quality. Compared to an MLP speculator without vertex-alignment, which suffers from quality loss, our method runs at a comparable speed.
5 Conclusion